An account was opened with $1,000 three years ago. Today, the account balance is $1,150. If

the interest is compounded annually, how much longer will it take for the account to have

earned a total of $300 in interest?
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

$1,150 = $1,000 * (1 + r)3

1.15 = (1 + r)3

1.04769 = 1 + r

r = 0.04769

r = 4.769%

When the interest accumulated = $3000, FV = $1,300

$1,300 = $1,000 * (1 + 0.04769)n

1.3 = 1.04769n

Taking log of both sides,

LN(1.3) = n * LN(1.04769)

0.2624 = n * 0.0466

n = 5.63 years